# From $r$ to $Q^{*}$: Your Language Model is Secretly a Q-Function

### v1

@@ Abstract
Reinforcement Learning From Human Feedback (RLHF) has been a critical to the success of the latest generation of generative AI models. In response to the complex nature of the classical RLHF pipeline, direct alignment algorithms such as Direct Preference Optimization (DPO) have emerged as an alternative approach. Although DPO solves the same objective as the standard RLHF setup, there is a mismatch between the two approaches. Standard RLHF deploys reinforcement learning in a specific token-level MDP, while DPO is derived as a bandit problem in which the whole response of the model is treated as a single arm. In this work we rectify this difference, first we theoretically show that we can derive DPO in the token-level MDP as a general inverse Q-learning algorithm, which satisfies the Bellman equation. Using our theoretical results, we provide three concrete empirical insights. First, we show that because of its token level interpretation, DPO is able to perform some type of credit assignment. Next, we prove that under the token level formulation, classical search-based algorithms, such as MCTS, which have recently been applied to the language generation space, are equivalent to likelihood-based search on a DPO policy. Empirically we show that a simple beam search yields meaningful improvement over the base DPO policy. Finally, we show how the choice of reference policy causes implicit rewards to decline during training. We conclude by discussing applications of our work, including information elicitation in multi-tun dialogue, reasoning, agentic applications and end-to-end training of multi-model systems.
@@ 1 Introduction
Empirically, we use our theoretical derivations to justify three practical insights which we believe to be of use to the community. First, we show that despite being derived as a contextual bandit, the implicit rewards of a DPO model have a per-token interpretation. Second, we demonstrate that likelihood search over a DPO model is analogous to searching over a reward function during decoding as done by contemporary works. Finally, we identify the choice of initial policy and reference distribution as being important in determining the trajectory of implicit rewards during training.
@@ 2 Related Work
Our work builds on foundational knowledge in maximum entropy RL (Ziebart, 2010) and inverse RL (Ziebart et al., 2008; Ng et al., 1999). In particular, we leverage the mapping between $Q$-functions and reward functions under a fixed policy as first done in inverse RL by Garg et al. (2022). Highly related to our work, Hejna & Sadigh (2024) exploit this relationship for RLHF, but for control in the traditional discounted MDP setting and Cundy & Ermon (2023) consider how LLMs can represent $Q$-functions for imitation learning.
@@ 4.1 DPO as a $Q$-function in the Token Level MDP
[…]
@@
In Ng et al. (1999)’s seminal work, the authors proved that two equivalent reward functions defined per definition 1 have the same optimal policy. By log-linearizing the optimal policy fixed point in eq. 5 and substituting in the Bellman equation from eq. 7, we have
@@
This is precisely the optimal advantage function, where $V^{*}$ directly follows the form of a potential shaping function. As also noted by contemporary works, using the advantage as reward preserves the optimal policy (Knox et al., 2024; Hejna et al., 2024). Unlike prior works, however, we demonstrate that this re-parameterization also leads to the same exact preference distribution as $r$.
@@
$r(\mathbf{s},\mathbf{a})=\beta\log\pi(\mathbf{s}|\mathbf{a})-\beta\log\pi_{\text{ref}}(\mathbf{s}|\mathbf{a})$ (11)
@@
Above we derived the invariance of the optimal policy under the re-parameterization. The preference model can be shown to be invariant by following the same steps used to arrive at eq. 8 in the last section. ∎
@@ 5.1 Does DPO Learn Credit Assignment?
In the previous section we outlined how the trained DPO policy represents an optimal Q-function for some reward that optimizes the preference equation. In this section, we evaluate qualitatively if the DPO-trained model is able to learn credit assignment from trajectory feedback. We begin with a generic set of Reddit posts for the TL;DR test dataset, which we provide in Appendix C with additional examples. In our representative example the user discusses an employment negotiations situation. Two answers are shown in Figure 1. The base summary, which is correct is provided on the left. On the right we modify the summary by introducing a higher-level position and a corresponding higher salary. For each token in both answers we compute the DPO-equivalent reward (advantage), $r(\mathbf{s},\mathbf{a})=\beta\log\pi_{\theta}(\mathbf{s}|\mathbf{a})-\beta\log\pi_{\text{ref}}(\mathbf{s}|\mathbf{a})$, where $\pi_{\theta}$ as outlined in Theorem 1 (here $\pi_{\theta}$ is our DPO-trained model and $\pi_{\text{ref}}$ is the SFT model). In Figure 1 each token is colored proportionally to this reward. We see that the model successfully identifies the tokens corresponding to the erroneous statements, while still maintaining comparable values for the rest, which is indicates that it can do credit assignment. Moreover, we see that within the context of the first error ("250K" salary) the model still allocates reasonable values to the rest of the tokens and specifically identifies the second error "management position". This is a promising sign of the ability to do "stitching" Levine et al. (2020). i.e. a form of combinatorial generalization from offline data. If this is the case, our findings could be significant for the use of reinforcement learning and RLHF in LLMs. At the same time, in the recently introduced RewardBench Lambert et al. (2024), DPO models have demonstrated strong performance as classifiers on reasoning tasks. We believe these are encouraging results, which warrant further large-scale study.
@@ 5.2 Connecting Guided Decoding and Search to Likelihood-Based DPO Optimization
where $\pi^{*}$ is the corresponding optimal policy. Now, since the starting state is fixed (it’s given by the prompt) we have that a search algorithm based on the conservative reward function of the RLHF objective and the corresponding optimal value policy is equivalent to likelihood search on the corresponding optimal policy. We empirically verify this property in Fig. 2, which shows the win rate of DPO models trained with three different $\beta$ values against the preferred summary in the test dataset. We see that a 5-beam search improves win-rates by 10-15% over the base policy (1-beam), which is comparable to the value-function guided search improvements reported in Mudgal et al. (2024). Interestingly, we see performance degrade with higher number of beams. Increasing the number of beams also produces answer with exploding length, which is a sign of reward over-optimization Gao et al. (2023); Park et al. (2024) and would explain the degradation in performance. These observations are consistent with out formulation of beam search as a search over a learned reward function.
@@
5.3 Can MaxEnt RL Explain DPO the Phenomenon Of Decreasing Likelihoods?
@@
While choosing $\pi_{\text{ref}}=\pi_{\text{ref}}^{w}$ is done in practice (Rafailov et al., 2023), it does mean that DPO will decrease the likelihood of all data in favor of extrapolated responses, which could cause over-fitting. Moreover, now that we have provided a derivation of DPO in the token-level MDP, one might expect it to exhibit characteristics like an RL algorithm – namely that the implied $Q$-function monotonically increases over time. However, this is not necesarily the case. Note that per analysis in Section section 3.1, DPO can be viewed as adjusting the reward (or advantage) function from which the optimal policy is deterministically mapped within the token-level MDP. DPO does not train a policy to maximize reward, and thus we do not argue about whether its implied value functions should increase or decrease over time.
@@ 6 Discussion
Learning reasoning from feedback: The STaR algorithm, Zelikman et al. (2022) improves reasoning in language models by sampling multiple reasoning chains and iteratively-fine-tuning on the ones that produce correct answers. Within the RLHF literature this is known as expert iteration and could produce sub-optimal performance as compared to DPO and PPO Dubois et al. (2024). If we formulate the reasoning problem as a ranking over generated answers, and based on Theorem 1, and our preliminary credit assignment results, it is promising that we can improve STaR by utilizing the filtered data as dis-preferred trajectories.
Multi-turn conversations: Teaching language models to be an interactive conversationalists has been difficult, as RLHF is optimized as a single-turn bandit formulation. Moreover, classical methods, such as PPO, are not applicable, since it is not generally possible to have human in the loop conversational training. Recent work by Andukuri et al. (2024) has shown success in this domain using STaR and extending DPO to multi-turn conversational trees such as the OpenAssistant dataset (Köpf et al., 2023) is a promising direction.
Agentic LLMs: LLM agents, such as WebGPT (Nakano et al., 2022) are equipped to take autonomous actions, such as browsing the Web and collecting information before providing an answer. The user then provides feedback based on the final output (in the original work this is still represented as pairwise preferences). Our derivations indicate that DPO training (on the full model trajectories) could learn optimal exploration behaviour. In a promising early work Song et al. (2024) show that iterative training using only sparse trajectory-level feedback, with DPO can improve significantly the success rate of an LLM agent on a number of tasks, even surpassing the performance of GPT4 on unseen environments.
End-to-end training of generative AI systems: Modern image generation systems, such as Dalle 3 Betker et al. (2023) use an LLM as a prompt refiner/generator in order to produce high quality conditioning before calling a diffusion image-generation model. If we consider a hybrid MDP with discrete token-level transitions, which generate a conditioning prompt that is the initial state for a separate diffusion MDP, this system can be optimized end-to-end with DPO. I.e. simultaneously improve the prompt drafter and image generation directly from user feedback. A potential example of such pipeline is included in Appendix D. Also, recent long-form video generation models Hu et al. (2023); Gupta et al. (2023) combine transformer-based auto-regressive generations with a diffusion-based decoder. The framework presented in Appendix D can potentially be applied to an RLHF pipeline in this setting as well.
We believe these are promising directions and plan to pursue them in further works.
@@ D.1 Hybrid Video Generative Models
Appendix E TL;DR Sample Generations
E.1 Samples Across Different Beam Searches
@@
E.2 Samples from DPO with and without SFT Pre-Training
@@
Appendix F Dataset and Hyperparameter Details
@@
Each TL;DR model was a pre-trained Pythia 2.8B model. SFT, when performed, was run for 1 epoch with batch size 128 and 8 gradient accumulation steps on 4 NVIDIA A40 GPUs. All DPO models were trained for 1 epoch with these same hyperparameters, along with a learning rate of $0.5\times 10^{-6}$ and linear warmup of 150 steps.

### v2

@@ Abstract
Reinforcement Learning From Human Feedback (RLHF) has been critical to the success of the latest generation of generative AI models. In response to the complex nature of the classical RLHF pipeline, direct alignment algorithms such as Direct Preference Optimization (DPO) have emerged as an alternative approach. Although DPO solves the same objective as the standard RLHF setup, there is a mismatch between the two approaches. Standard RLHF deploys reinforcement learning in a specific token-level MDP, while DPO is derived as a bandit problem in which the whole response of the model is treated as a single arm. In this work we rectify this difference. We theoretically show that we can derive DPO in the token-level MDP as a general inverse Q-learning algorithm, which satisfies the Bellman equation. Using our theoretical results, we provide three concrete empirical insights. First, we show that because of its token level interpretation, DPO is able to perform some type of credit assignment. Next, we prove that under the token level formulation, classical search-based algorithms, such as MCTS, which have recently been applied to the language generation space, are equivalent to likelihood-based search on a DPO policy. Empirically we show that a simple beam search yields meaningful improvement over the base DPO policy. Finally, we show how the choice of reference policy causes implicit rewards to decline during training. We conclude by discussing applications of our work, including information elicitation in multi-turn dialogue, reasoning, agentic applications and end-to-end training of multi-model systems.
@@ 1 Introduction
Empirically, we use our theoretical derivations to justify three practical insights which we believe to be of use to the community. First, we show that despite being derived as a contextual bandit, the implicit rewards of a DPO model have a per-token interpretation. Second, we demonstrate that likelihood search over a DPO model is analogous to searching over a reward function during decoding as done by contemporary works (Liu et al., 2023b; Feng et al., 2024). Finally, we identify the choice of initial policy and reference distribution as being important in determining the trajectory of implicit rewards during training.
@@ 2 Related Work
Our work builds on foundational knowledge in maximum entropy RL (Ziebart, 2010) and inverse RL (Ziebart et al., 2008; Ng et al., 1999; Cao et al., 2021). In particular, we leverage the mapping between $Q$-functions and reward functions under a fixed policy as first done in inverse RL by Garg et al. (2022). Related to our work, Nachum et al. (2017) uses similar derivations for reinforcement learning in control and Watson et al. (2023) does so for inverse RL from demonstration. Hejna & Sadigh (2024) exploit this relationship for RLHF. While related, these works still require an additional loop of reinforcement learning optimization, which we dispose of in our formulation of feedback learning for LLMs. In the LLM domain, Yu et al. (2024) uses pre-trained models as priors for Q-learning, while Cundy & Ermon (2023) considers a similar formulation for imitation learning. In this work instead, we formulate preference-based learning as Q-learning.
@@ 4.1 DPO as a $Q$-function in the Token Level MDP
$V^{*}(\mathbf{s}_{t})=\beta\log\sum_{\mathbf{a}\in\mathcal{A}}e^{Q^{*}(\mathbf{s}_{t},\mathbf{a})/\beta}$ (6)
@@
In Ng et al. (1999)’s seminal work, the authors proved that two equivalent reward functions defined per definition 1 have the same optimal policy. By log-linearizing the optimal policy fixed point in eq. 5 and substituting in the Bellman equation from eq. 7 (Nachum et al., 2017; Watson et al., 2023), we have
@@
This is precisely the optimal advantage function, where $V^{*}$ directly follows the form of a potential shaping function. Watson et al. (2023) first used this derivation to arrive at a “coherent” reward function and follow-ups arrived at the same conclusion by noting that using the advantage as reward preserves the optimal policy (Knox et al., 2024; Hejna et al., 2024). Unlike prior works, however, we demonstrate that this re-parameterization also leads to the same exact preference distribution as $r$.
@@
$r(\mathbf{s},\mathbf{a})=\beta\log\pi(\mathbf{a}|\mathbf{s})-\beta\log\pi_{\text{ref}}(\mathbf{a}|\mathbf{s})$ (11)
@@
Above we derived the invariance of the optimal policy under the re-parameterization. The preference model can be shown to be invariant by substituting and following the same steps used to arrive at eq. 8 in the last section, or by following Definition 1 from Watson et al. (2023). ∎
@@ 5.1 Does DPO Learn Credit Assignment?
In the previous section we outlined how the trained DPO policy represents an optimal Q-function for some reward that optimizes the preference equation. In this section, we evaluate qualitatively if the DPO-trained model is able to learn credit assignment from trajectory feedback. We begin with a generic set of Reddit posts for the TL;DR test dataset, which we provide in Appendix C with additional examples. In our representative example the user discusses an employment negotiations situation. Two answers are shown in Figure 1. The base summary, which is correct is provided on the left. On the right we modify the summary by introducing a higher-level position and a corresponding higher salary. For each token in both answers we compute the DPO reward (equivalently the advantage function or “coherent” reward (Watson et al., 2023)), $r(\mathbf{s},\mathbf{a})=\beta\log\pi_{\theta}(\mathbf{s}|\mathbf{a})-\beta\log\pi_{\text{ref}}(\mathbf{s}|\mathbf{a})$, where $\pi_{\theta}$ as outlined in Theorem 1 (here $\pi_{\theta}$ is our DPO-trained model and $\pi_{\text{ref}}$ is the SFT model). In Figure 1 each token is colored proportionally to this reward. We see that the model successfully identifies the tokens corresponding to the erroneous statements, while still maintaining comparable values for the rest, which is indicates that it can do credit assignment. Moreover, we see that within the context of the first error ("250K" salary) the model still allocates reasonable values to the rest of the tokens and specifically identifies the second error "management position". This is a promising sign of the ability to do "stitching" Levine et al. (2020) i.e. a form of combinatorial generalization from offline data. If this is the case, our findings could be significant for the use of reinforcement learning and RLHF in LLMs, particularly for compositional tasks, such as code and reasoning. At the same time, in the recently introduced RewardBench Lambert et al. (2024), DPO models have demonstrated strong performance as classifiers on reasoning tasks. We believe these are encouraging results, which warrant further large-scale study beyond our qualitative observations.
@@ 5.2 Connecting Guided Decoding and Search to Likelihood-Based DPO Optimization
where $\pi^{*}$ is the corresponding optimal policy. Now, since the starting state is fixed (it’s given by the prompt) we have that a search algorithm based on the conservative reward function of the RLHF objective and the corresponding optimal value policy is equivalent to likelihood search on the corresponding optimal policy. We empirically verify this property in Fig. 2, which shows the win rate of DPO models trained with three different $\beta$ values against the preferred summary in the test dataset. We see that a 5-beam search improves win-rates by 10-15% over the base policy (1-beam), which is comparable to the value-function guided search improvements reported in Mudgal et al. (2024). Interestingly, we see performance degrade with higher number of beams. Increasing the number of beams also produces answer with exploding length, which is a sign of reward over-optimization Gao et al. (2023); Park et al. (2024); Rafailov et al. (2024) and would explain the degradation in performance. These observations are consistent with out formulation of beam search as a search over a learned reward function.
@@
5.3 Connections Between Proxy Tuning and Reinforcement Learning
Several recent works Mitchell et al. (2023); Liu et al. (2024a; b) have proposed an approach of inference-time model alignment through a proxy guidance model. These approaches start with a (unaligned) base model $\pi_{\text{base}}$ and a proxy model $\pi_{\text{proxy}}$ and a target distribution reference model $\pi_{\text{ref}}$. The inference time re-alignment of the base model is carried by re-weighting the conditional probabilities of each token:
$\pi(\mathbf{a}|\mathbf{s}_{t})\propto\pi_{\text{base}}(\mathbf{a}|\mathbf{s}_{t})\left(\frac{\pi_{\text{proxy}}(\mathbf{a}|\mathbf{s}_{t})}{\pi_{\text{ref}}(\mathbf{a}|\mathbf{s}_{t})}\right)^{\beta}$ (14)
Under our considerations from the prior chapter, then this becomes equivalent to
$\pi(\mathbf{a}|\mathbf{s}_{t})\propto\pi_{\text{base}}(\mathbf{a}|\mathbf{s}_{t})\exp(\beta(Q^{*}(\mathbf{s}_{t},\mathbf{a})-V^{*}(\mathbf{s}_{t})))$ (15)
where $\beta(Q^{*}(\mathbf{s}_{t},\mathbf{a})-V^{*}(\mathbf{s}_{t})$ is the optimal implicit advantage from the proxy tuning model. That is our theoretical results allows us to tie the realignment approaches of Mitchell et al. (2023); Liu et al. (2024a; b) to recent works which explicitly train critic models Mudgal et al. (2024) for token-level decoding.
5.4 Likelihoods should decrease when using DPO.
@@ 5.3 Connections Between Proxy Tuning and Reinforcement Learning
While choosing $\pi_{\text{ref}}=\pi_{\text{ref}}^{w}$ is done in practice (Rafailov et al., 2023), it does mean that DPO will decrease the likelihood of all data in favor of extrapolated responses, which could cause over-fitting. Moreover, now that we have provided a derivation of DPO in the token-level MDP, one might expect it to exhibit characteristics like an RL algorithm – namely that the implied $Q$-function monotonically increases over time. However, this is not necesarily the case. Note that per analysis in Section section 3.1, DPO can be viewed as adjusting the reward (or advantage) from which the optimal policy is deterministically mapped within the token-level MDP. DPO does not train a policy to maximize reward, and thus we do not argue about whether its implied value functions should increase or decrease over time.
@@ 6 Discussion
Learning intermediate reasoning from outcome feedback: Recent works have shown promising results on that front Pang et al. (2024); Hwang et al. (2024).
Multi-turn conversations: Teaching language models to be an interactive conversationalists has been difficult, as RLHF is optimized as a single-turn bandit formulation. Moreover, classical methods, such as PPO, are not applicable in this setting. Recent work by Andukuri et al. (2024) has shown success in this domain using STaR and extending DPO to multi-turn conversational trees is a promising direction.
Agentic LLMs: LLM agents, such as WebGPT (Nakano et al., 2022) are equipped to take autonomous actions, such as browsing the Web and collecting information before providing an answer. The user then provides feedback based on the final output. Our derivations indicate that DPO training (on the full model trajectories) could learn optimal exploration behaviour. Recent works Song et al. (2024); Xi et al. (2024) shows promise in that direction.
End-to-end training of generative AI systems: Modern image generation systems, such as Dalle 3 Betker et al. (2023) use an LLM to produce high quality conditioning before calling a diffusion generation model. Also, recent long-form video generation models Hu et al. (2023); Gupta et al. (2023) combine transformer-based auto-regressive generations with a diffusion-based decoder. Such systems could potentially be optimized end-to-end with ahybrid version of DPO. We expand on these points in the Appendix.
We believe these are promising directions for future work.
@@ D.1 Hybrid Video Generative Models
Appendix E Beam Search Trends for PPO
We consider whether the beam search trends in Figure 2 hold for PPO, since PPO is known to exhibit reward over-optimization as well (Gao et al. (2023)). We use a PPO-tuned GPT-J-6B model released by CarperAI [^1] https://huggingface.co/CarperAI/openai_summarize_tldr_ppo, fine-tuned on the same TL;DR dataset. We use the same sampling parameters as with the DPO experiments ($\tau=1.0$, $k=50$) and generate $256$ samples from test set prompts, with GPT-4 as the evaluator. We only report results for 1, 2, 5, and 7 beams; higher number of beams were tried but exhausted memory on an NVIDIA A40 GPU.
In Figure 5, we observe a similar over-optimization phenomenon as with DPO, with 2 beams both increasing winrate and decreasing sample length (even under the length-biased evaluator GPT-4). However, more than 2 beams leads to a decline in downstream performance and large uptick in sample length, similar behavior to DPO in the over-optimization regime. We find that the benefits of increasing the beam size are more limited when using this model.
Figure 5: PPO model performance using beam search. Left: Win-rate of model summaries over preferred summary. Right: Average answer length based on number of beams.
Appendix F TL;DR Sample Generations
F.1 Samples Across Different Beam Searches
@@
F.2 Samples from DPO with and without SFT Pre-Training
@@
Appendix G Dataset and Hyperparameter Details
@@
Each TL;DR model was a pre-trained Pythia 2.8B model. We first SFT with a learning rate of $0.5\times 10^{-6}$ and batch size of 128 for one epoch using 8 gradient accumulation steps on 4 NVIDIA A40 GPUs. All DPO models with the same setup and hyperparameters as the original paper (Rafailov et al., 2023). Specifically, we train for 1 epoch with a learning rate of $0.5\times 10^{-6}$ and linear warmup of 150 steps. We use a batch size of 32 (example example is a positive and a negative) and clip gradient norms to 10. We use $\beta=0.1$ for DPO.